OpenStudy (anash):

A rectangular national hockey league ice rink has a perimeter 570 ft. The length of the rink is 30 ft longer than twice the width. What are the dimensions of an NHL ice rink?

OpenStudy (shamil98):

p = 2l + 2w L = 30 + 2w 570 = 2(30+2w) + 2w 570 = 60 + 4w + 2w 510 = 6w 85 = w L = 30 + 2(85) L = 30 + 170 L = 200